OpenStudy (anonymous):

Find the probability of the event when a number cube with faces labeled 1– 6 is rolled once. a number greater than 4 A. 2/3 B.1/3 C.1/6 D.1/2

OpenStudy (gunsodo):

N(E) Number of events with following condition = {5, 6} = 2 N(S) Number of events with no condition = {1, 2, 3, 4, 5, 6} = 6 P(E) = N(E)/N(S) = 2/6 = 1/3 Ans : B